OpenStudy (anonymous):

Solve each radical equation by using algebra. Check your solution. I will use sq. rt for square root. 1. sq. rt of (x- 5) = 3 2. sq. rt of ( 2x - 1) = x 3. 3 sq. rt of( x - 2) = 2 sq. rt of( x) 4. sq. rt. of (3x - 2) = x - 2

OpenStudy (anonymous):

PLEASE HELP THIS IS DUE IN 4 HOURS

OpenStudy (anonymous):

1. \[\Large \sqrt{2x-1}=x\] is the first one like this??

OpenStudy (anonymous):

second one..

OpenStudy (anonymous):

Thats exactly how its written, exactly how i typed it

OpenStudy (anonymous):

but you said you'll use sq.rt for square root, then why did you type sq.rt before equation ? :S ... ? could you draw them please (equations)

OpenStudy (anonymous):

1. \[(\sqrt{x-5})^{2}=(3)^2\] x-5=9 x=14

OpenStudy (anonymous):

2. \[(\sqrt{2x-1})^{2}=(x)^{2}\] 2x-1=x^2 -x^2+2x-1 --> (solve it with Quadratic quation to find x)

OpenStudy (anonymous):

3. \[(3\sqrt{x-2})^{2}=(2\sqrt{x})^{2}\] 9(x-2)=4(x) 9x-18=4x 9x-4x=18 5x=18 x=3.6

OpenStudy (anonymous):

4. \[(\sqrt{3x-2})^{2}=(x-2)^{2}\] \[3x-2=x ^{2}-4x+4\] 7x-6=x^2 -x^2+7x-6 -> (solve with quadratic equation)